OpenStudy (anonymous):

Evaluate the expression and write the result in the form a + bi. (−5 + i) − (8 − 5i)

OpenStudy (anonymous):

-5+i-8+5i=-13+6i

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can it not go further

OpenStudy (anonymous):

no the question said write the result in the form a + bi. =-13+6i

OpenStudy (anonymous):

o i see thanks again